OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!!!!!!! A sandbag was thrown downward from a building. The function f(t) = -16t^2 - 64t + 512 shows the height f(t), in feet, of the sandbag after t seconds. Complete the square of the expression for f(x) to determine the vertex of the graph of f(x). Would this be a maximum or minimum on the graph?

OpenStudy (anonymous):

@Hero @ganeshie8 @tHe_FiZiCx99 @zzr0ck3r

ganeshie8 (ganeshie8):

where are you stuck ?

ganeshie8 (ganeshie8):

f(t) = -16t^2 - 64t + 512 you need to complete the square, right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

im so bad at it

ganeshie8 (ganeshie8):

basically we need to change it to this form : \[\large f(c) = a(x-h)^2 + k\]

ganeshie8 (ganeshie8):

start by factoring out "-16" from first two terms : \[\large \begin{align} \\ f(t) &= -16t^2 - 64t + 512 \\ &= -16(t^2+4t) + 512\end{align}\]

ganeshie8 (ganeshie8):

next, recall the identity : \(\large \color{red}{(a+b)^2 = a^2+2ab+b^2}\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it will be a maximum, as the graph gets more negative as t increases

OpenStudy (anonymous):

so it turns into\[f(t) = -16(t^2 + 4t + 4) + 512\]

OpenStudy (anonymous):

right?

ganeshie8 (ganeshie8):

Perfect !

OpenStudy (anonymous):

after factorising you get -16[(t+2)^2 - 36]

ganeshie8 (ganeshie8):

wait a sec, you have added 4, so u better subtract it also, right ?

OpenStudy (anonymous):

right :)

ganeshie8 (ganeshie8):

the stuff inside parenthesis can be written into this form, if we're clever enough we will see it immediately : start by factoring out "-16" from first two terms : \[\large \begin{align} \\ f(t) &= -16t^2 - 64t + 512 \\ &= -16(t^2+4t) + 512 \\ &= -16(t^2+4t + \color{red}{2^2} ) + 512 - \color{red}{(-16*2^2)}\end{align}\]

ganeshie8 (ganeshie8):

see if that looks okay ^^

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

so then u simplify?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

see if you get -16[(t+2)^2 - 36]

OpenStudy (anonymous):

after that you can sub in t= 1 and t=2 to check if it is a maximum or minimum curve

ganeshie8 (ganeshie8):

yes before simplifying, complete the square so that we feel accomplished a bit :) \[\large \begin{align} \\ f(t) &= -16t^2 - 64t + 512 \\ &= -16(t^2+4t) + 512 \\ &= -16(t^2+4t + \color{red}{2^2} ) + 512 - \color{red}{(-16*2^2)} \\ &= -16(t+\color{red}{2})^2 + 512 - \color{red}{(-16*2^2)} \\ \end{align}\]

OpenStudy (anonymous):

i got\[f(x) = -16(t+2)^2 + 576\]

OpenStudy (anonymous):

am i right @ganeshie8 ?

ganeshie8 (ganeshie8):

Looks good ^^

ganeshie8 (ganeshie8):

compare it with : \[\large f(x) = a(x-\color{Red}{h})^2 + \color{Red}{k}\] vertex = \((\color{Red}{h}, \color{Red}{k})\) = ?

OpenStudy (anonymous):

(-2, 576) right?

ganeshie8 (ganeshie8):

Correct !!

ganeshie8 (ganeshie8):

so the value of function at vertex is 576 Is that a maximum or minimum value ?

OpenStudy (anonymous):

maximum, right?

ganeshie8 (ganeshie8):

thats right ! but may i knw why do you think it is maximum ? :)

OpenStudy (anonymous):

cuz the leading coefficient of the original equation is negative. :)

ganeshie8 (ganeshie8):

Excellent ! good job :)

OpenStudy (anonymous):

THANK YOU SO MUCH!!!!!! :) <3 :) <3 ✱